OpenStudy (anonymous):

could anyone help me step by step on how to do this? find the value of the variable [-12 -w^2] = [2k -81] [2f 3 ] [-14 3] 2. evaluate the determinant of the matrix {-4 5 6] [0 4 4] [-2 -5 4]

OpenStudy (anonymous):

[-12 -w^2] = [2k -81] comparing the terms at the like places we have: -12= 2k k=-6 -w^2= -81 w^2= 81 w = 9

OpenStudy (anonymous):

[-12 -w^2] = [2k -81] [2f 3 ] = [-14 3] comparing the terms at the like places we have: -12= 2k k=-6 -w^2= -81 w^2= 81 w = 9 2f = -14 f= -7

OpenStudy (rock_mit182):